At the beginning of his current tax year, Eric bought a corporate bond with a maturity value of $32,000 from the secondary market for $24,500. The bond has a stated annual interest rate of 7 percent payable on June 30 and December 31, and it matures in five years on December 31.

Absent any special tax elections, how much interest income will Eric report from the bond this year and in the year the bond matures?


Interest income reported this year?

Interest income reported in the year the bond matures?
Answer #1

INTEREST INCOME REPORTED THIS YEAR :

On June 30 = Maturity value × 7% × (6÷12) = $32000 × 7% × (6÷12) = $1120.

On Dec. 31 = Maturity value × 7% × (6÷12) = $32000 × 7% × ( 6÷12) = $1120.

Total interest income reported this year = $1120 + $1120 = $2240.

INTEREST INCOME REPORTED IN THE YEAR BOND MATURES :

= Interest income related to market discount + interest income of the year.

= ( $32000 - $24500 ) + $2240

= $7500 + $2240

= $9740.

______×______

ALL THE BEST

Keep Learning